# Group inverses of matrices of simple symmetric directed graphs

### v1

@@ 3. Necessary and Sufficient Condition of $A$ such that $D(A),D(A^{\#})\in\mathbb{D}$
Since the zero non-zero pattern of $A$ and $P^{-1}A^{\#}P$ are same, $D(A^{\#})$ is also a corona graph. ∎
Finally, we classify all the real square matrix $A$ such that $D(A)$ and $D(A^{\#})$ both are in the class $\mathbb{D}$, that is $D(A)$ is either a corona graph or a star graph.
Theorem 3.5.
Let $A$ be an $n\times n$ real matrix such that $D(A)\in\mathbb{D}$. Assume that $D(A)$ is strongly connected and $\Delta_{A}\neq 0$. Then $D(A^{\#})\in\mathbb{D}$ if and only if $D(A)$ is either a corona graph or a star tree graph.
Proof.
Let $D(A)$ be a corona or star tree graph. Then, by Proposition 3.4, $D(A^{\#})\in\mathbb{D}$. Conversely, suppose $D(A)$ is neither a corona nor a star tree graph. Since $D(A)\in\mathbb{D}$, there exists a non-pendant vertex adjacent to more than one pendant vertices and at least two non-pendant vertices in $D(A)$. Let $i$ be a non-pendant vertex adjacent to $s\geq 2$ pendant vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ and $i$ is adjacent to another non-pendant vertex $j$. The alternating paths that start or end with vertex $i$ are $((i,i_{m},i))$ for all $1\leq m\leq s$. So, $i$ is adjacent to exactly $s$ vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ in $D(A^{\#})$. Let $j$ be adjacent to a pendant vertex $\tilde{j}$. Then, for any $1\leq m\leq s$, $((i_{m},i,i_{m}))$ and $((i_{m},i,i_{m}),(i,j,i),(j,\tilde{j},j))$ are two alternating paths start from vertex $i_{m}$. So, vertex $i_{m}$ is non-pendant in $D(A^{\#})$ for all $1\leq m\leq s$. Since we have a non-pendant vertex $i$ which is not adjacent to any pendant vertices in $D(A^{\#})$, $D(A^{\#})$ does not belong to $\mathbb{D}$. ∎
@@ Acknowledgments
The author was supported by Indian Institute of Technology Gandhinagar Post-Doctoral Fellowship IP/IITGN/MATH/PC/2223/20.

### v3

@@ 3. Necessary and sufficient condition of a real matrix $A$ such that $D(A),D(A^{\#})\in\mathcal{D}$
Since the zero non-zero pattern of $A$ and $P^{-1}A^{\#}P$ are same, $D(A^{\#})$ is also a corona digraph. ∎
Finally, we classify all the real square matrix $A$ such that $D(A)$ and $D(A^{\#})$ both are in the class $\mathcal{D}$, that is $D(A)$ is either a corona digraph or a star digraph.
Proof of Theorem 1.5.
Let $D(A)$ be a corona or star tree digraph. Then, by Proposition 3.5, $D(A^{\#})\in\mathcal{D}$. Conversely, suppose $D(A)$ is neither a corona nor a star tree digraph. Since $D(A)\in\mathcal{D}$, there exists a non-pendant vertex adjacent to more than one pendant vertices and at least two non-pendant vertices in $D(A)$. Let $i$ be a non-pendant vertex adjacent to $s\geq 2$ pendant vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ and $i$ is adjacent to another non-pendant vertex $j$. The alternating cycle chains that start or end with vertex $i$ are $((i,i_{m},i))$ for all $1\leq m\leq s$. So, $i$ is adjacent to exactly $s$ vertices $\{i_{1},i_{2},\ldots,i_{s}\}$ in $D(A^{\#})$. Let $j$ be adjacent to a pendant vertex $\tilde{j}$. Then, for any $1\leq m\leq s$, $((i_{m},i,i_{m}))$ and $((i_{m},i,i_{m}),(i,j,i),(j,\tilde{j},j))$ are two alternating cycle chains start from vertex $i_{m}$. So, vertex $i_{m}$ is non-pendant in $D(A^{\#})$ for all $1\leq m\leq s$. Since we have a non-pendant vertex $i$ which is not adjacent to any pendant vertices in $D(A^{\#})$, $D(A^{\#})$ does not belong to $\mathcal{D}$. ∎
@@ Acknowledgments
The author thanks the Indian Institute of Information Technology, Design & Manufacturing, Kurnool for providing a good research environment.
Declarations
Conflict of Interest: The authors declare that there is no conflict of interest with anyone.